OpenStudy (anonymous):

Solve the equation. 3x - 3= -18 x = -7 x = -5 x = -3 x = 5 I will medal!!

OpenStudy (greatlife44):

You need to isolate x do you know what to do @rainbowswirl

OpenStudy (anonymous):

Not exactly @greatlife44

OpenStudy (greatlife44):

Okay you've got 3x-3 = -18 right? we have 3x -3 this is subtraction. we must do the opposite of subtraction what's that?

OpenStudy (anonymous):

addition @greatlife44

OpenStudy (greatlife44):

Yes. we and we must do this to both sides of our equation so what would that be

OpenStudy (anonymous):

What do I add to both sides? @greatlife44

OpenStudy (greatlife44):

3 but can you see why based on what I just said?

OpenStudy (anonymous):

Oh, yes, now I see it. So, add 3 to both sides of 3x -3?? @greatlife44

OpenStudy (greatlife44):

excellent

OpenStudy (greatlife44):

Continue and write out what we have so far

OpenStudy (anonymous):

Isn't that 6 on each side?? @greatlife44

OpenStudy (greatlife44):

3x-3 = -18 +3 +3 3x = -15

OpenStudy (greatlife44):

Notice how when we add 3 to both sides 3 cancels out because -3+3 = 0 so we're just left with 3x = -18+3

OpenStudy (anonymous):

Okay, what do you do next? @greatlife44

OpenStudy (greatlife44):

we have 3x = -15

OpenStudy (greatlife44):

what do you think we need to do now?

OpenStudy (anonymous):

B. x=-5 source: http://www.wolframalpha.com/input/?i=3x-3%3D-18

OpenStudy (anonymous):

I'm not sure, that why im struggling @greatlife44

OpenStudy (anonymous):

thats*

OpenStudy (study312):

-18 + 3 = -15. -15 ÷ 3 = -5. x = -5

OpenStudy (greatlife44):

giving him a direct answer won't help him

OpenStudy (study312):

sorry, you guys already did that so I was just trying to show it was right

OpenStudy (greatlife44):

her*

OpenStudy (greatlife44):

think about it we have 3x = -15 3x means three times x we need to do the opposite of multiplication to isolate x.

OpenStudy (anonymous):

Can anyone help me with more??

OpenStudy (misssmartiez):

3 x -5 = -15. -15 - 3 = 18

OpenStudy (misssmartiez):

sure.

OpenStudy (misssmartiez):

Remember, negative are opposed. So the less numbers in positive, would be the greater numbers, since close to zero.

OpenStudy (anonymous):

Ok^

OpenStudy (study312):

also if you take -3 and move it to the -18 you get -18 - 3 now because you brought over the - 3 you change it to + 3, get it? so it becomes x -3 + 3 = -18 + 3

OpenStudy (anonymous):

Solve the equation. r/2 -6 =14 r =40 r =34 r =20 r =16

OpenStudy (study312):

please close this post and make a new one with your new question

OpenStudy (misssmartiez):

40

OpenStudy (misssmartiez):

Do it yourself, 40/2 = what?

OpenStudy (misssmartiez):

then - that by 6. Actually easy, just think about it :)!

OpenStudy (misssmartiez):

I don't mind if you put it into this post.

OpenStudy (anonymous):

Ok @MissSmartiez

OpenStudy (misssmartiez):

Fire away.